# MoD-SLAM: Monocular Dense Mapping for Unbounded 3D Scene Reconstruction

### 21 Feb 2024 v3

@@ 4.2 Results and Evaluation
Evaluation on Replica [28] Table 1 shows the results of the reconstruction of systems on the Replica dataset. Compared to the systems mentioned in the table, our system exhibits superior tracking capabilities in both monocular and RGB-D modes. In comparison with NeRF-based SLAM [46, 43, 45], our approach achieves enhanced reconstruction fidelity. Compared to 3D Gaussian-based SLAM systems [11], our system falls in terms of scene reconstruction accuracy. However, our system outperforms in tracking performance and is more effective in recovering the scale of the scene. Fig. 4 presents a comparative analysis of scene reconstructions by our system against NICE-SLAM [46] and GO-SLAM [43]. Our method demonstrates superior performance that shows smoother scene reconstruction.
@@ 5 Conclusion
Limitations The depth estimation module relies on a pre-trained model; thus, for complex scenes or those significantly divergent from our provided pre-trained scenes, it’s necessary to retrain the depth estimation module. Moreover, compared to traditional SLAM, our method lacks an IMU [32] module, leading to tracking errors. Finally, the integration of depth estimation in mapping under monocular mode results in a slower mapping speed.
@@ 0.A.1 Ablation Study
0.A.2 Runtime Analysis
We test the tracking and the mapping time of NICER-SLAM [45], GO-SLAM [43] and our system using the same number of sampling points. Fig. 7(b) shows that our system is faster than GO-SLAM and NICER-SLAM in the tracking process. However, because of the depth estimation module, our system is slightly slower than GO-SLAM in the mapping process, but higher than NICER-SLAM. In addition, Fig. 7(b) also demonstrates that our model size is smaller and uses less memory. GO-SLAM and our system are benchmarked with an NVIDIA RTX 4090 GPU while NICER-SLAM is benchmarked with an A100.
Scene fr1-desk fr2-xyz fr3-office
GO-SLAM[43] 1.500 0.600 1.000
RGB-D Co-SLAM[38] 2.400 1.700 2.400
Ours 1.382 0.641 0.939
DROID-SLAM[34] 2.713 1.304 2.573
DeepV2D[33] 21.391 6.950 11.922
Mono ORB-SLAM3[4] 1.534 0.720 1.400
GO-SLAM[43] 2.381 1.093 1.773
Ours 1.481 0.686 1.101
Table 6: ATE RMSE [cm] on the TUM dataset [29]. The performance metrics reported by our system for each scene are derived from the average of three executions per scene.
0.A.3 Additional Experimental Results
TUM Dataset [29] We further test the tracking accuracy on the TUM dataset [29]. As shown in Table 6, Our method not only achieve the most accurate tracking performance in RGBD mode, but also outperforms classic SLAM systems as well as GO-SLAM [43] in monocular mode.
Replica Dataset [29] Here we show more results on Replica [29]. Fig. 8, Fig. 9 and Fig. 7(a) demonstrate the superior scene reconstruction capability of our proposed method in monocular and RGB-D mode, and also exhibiting enhanced smoothness in comparison to GO-SLAM [43].

### 26 Feb 2024 v4

@@ 4.2 Results and Evaluation
Evaluation on Replica [28] Table 1 shows the results of the reconstruction of systems on Replica. Compared to the systems mentioned in the table, our system exhibits superior tracking capabilities in both monocular and RGB-D modes. In comparison with NeRF-based SLAM [46, 43, 45], our approach achieves enhanced reconstruction fidelity. Compared to 3D Gaussian-based SLAM systems [11], our system falls in terms of scene reconstruction accuracy. However, our system outperforms in tracking performance and is more effective in recovering the scale of the scenes. Fig. 4 and Fig. 7(a) present a comparative analysis of scene reconstructions by our system against NICE-SLAM [46] and GO-SLAM [43]. Our method demonstrates superior performance.
@@
Evaluation on TUM-RGBD [29] To further test our method, we compared our method with point-cloud SLAM methods [34, 33, 4] and NeRF-based SLAM methods [43, 38] on the TUM RGB-D dataset. Since the TUM RGB-D dataset lacks ground truth meshes, we compared the accuracy of the tracking process. Table 3 shows the accurate tracking performance of our method.
Runtime Analysis We tested the tracking and the mapping time of NICER-SLAM [45], GO-SLAM [43] and our system using the same number of sampling points. Fig. 7(b) shows that our system is faster than GO-SLAM and NICER-SLAM in the tracking process. However, because of the depth estimation module, our system is slightly slower than GO-SLAM in the mapping process, but higher than NICER-SLAM. In addition, Fig. 7(b) also demonstrates that our model size is smaller and uses less memory.
@@ 5 Conclusion
Limitations The depth estimation module relies on a pre-trained model; thus, for complex scenes or those significantly divergent from our provided pre-trained scenes, it’s necessary to retrain the depth estimation module. Moreover, compared to traditional SLAM, our method lacks an IMU [32] module, leading to tracking errors.
@@ 0.A.1 Ablation Study
0.A.2 Additional Experimental Results
Replica Dataset [29] Here we show more results on Replica [29]. Fig. 8, Fig. 9 and demonstrate the superior scene reconstruction capability of our proposed method in monocular and RGB-D mode, and also exhibiting enhanced smoothness in comparison to GO-SLAM [43].